Tables at a wedding reception have six sides or eight sides and will be trimmed around the edges with lace. Each side on a given table is the same length, but a side on a six-sided table is 8 inches longer than a side on ane eight-sided table. One package of lace trims the eight-sided table with no excess. When the same package of lace trims the six-sided table, there are 2 inches of lace left over. What is the length of lace in one package
Answer by stanbon(48543)

Tables at a wedding reception have six sides or eight sides and will be trimmed around the edges with lace.
Each side on a given table is the same length, but a side on a six-sided table is 8 inches longer than a side on ane eight-sided table.
One package of lace trims the eight-sided table with no excess.
When the same package of lace trims the six-sided table, there are 2 inches of lace left over.
What is the length of lace in one package.
----------------------------------------------
8-sided table
Let length of side be "x" inches
Total length of lace = 8x
------------------------
6-sided table
length of side = "x+8" inches
Total length of lace = 6(x+8)+2
-----------------------------------
Equation:
8x = 6(x+8)+2
8x = 6x + 50
2x = 50
x = 25 inches
---------------------
Total length of lace = 8*25 = 200 inches.
================================================
Cheers,
Stan H.
---------